# PoLAR: Polar-Decomposed Low-Rank Adapter Representation

### v1

@@ F.3 Approximate Feasibility
F.4 Gemma-3-12B on Commonsense Reasoning
@@ G.4 Matrix Factorization
Here, we provide the experimental details for Fig. 3. We compare Riemannian Gradient Descent (RGD) with the PoLAR parameterization (PO) to Gradient Descent (GD) with the Burer-Monteiro parameterization (BM). We set $\mathbf{A}\in\mathbb{R}^{m\times n}$ with $r_{A}=4$ evenly spaced singular values in the interval $[1,\kappa]$ with $\kappa\in\{10,100\}$. For $\kappa=10$, we set $\eta=10^{-3}$ and $\gamma=1$ for PO + RGD and $\eta=10^{-3}$ for BM + GD. For $\kappa=100$, we use $\eta=10^{-4}$ for both methods. In the symmetric case, we use again $\eta=10^{-3}$ for $\kappa=10$ and $\eta=10^{-5}$ for $\kappa=100$.

### v2

@@ F.3 Approximate Feasibility
F.4 Comparison with GaLore
We compare PoLAR to GaLore (Zhao et al., 2024) in Table 8. We adopt our commonsense reasoning training setup from Table 2 with rank $16$. For GaLore’s hyperparameters, we take inspiration from their fine-tuning experiments in Appendix D and set the scale to $2$. We set the update projection gap s.t. the projections are updated once 15% increments of training are completed and tune the learning rate $\eta\in\{1\times 10^{-5},4\times 10^{-5}\}$.
Table 8: Performance on commonsense reasoning tasks with Llama-2-7B using GaLore and PoLAR for rank 16 in a single-task setup. HeSw refers to HellaSwag and WiGr to WinoGrande.
$\eta$ BoolQ PIQA SIQA HeSw WiGr ARC-e ARC-c OBQA Avg.
GaLore $1\times 10^{-5}$ 81.93 80.20 56.40 80.23 81.53 80.22 47.53 44.80 69.10
$4\times 10^{-5}$ 86.48 82.81 61.26 83.84 78.85 83.08 54.44 57.80 73.57
PoLAR 88.01 82.92 59.93 82.68 81.77 81.82 57.51 57.40 74.00
F.5 Gemma-3-12B on Commonsense Reasoning
@@ G.2 Mathematical Reasoning
For the mathematical reasoning experiment, we tune the learning rate grid in $\{2\times 10^{-4},4\times 10^{-4},6\times 10^{-4}\}$. We train for 2 epochs on MetaMathQA (Yu et al., 2024) using rank $16$, batch size $128$, and tune $\lambda\in\{10^{-3},5\times 10^{-3}\}$. Further following Yu et al. (2024), we do not use few-shot prompting, as the zero-shot setup reduces inference costs and tends to work better for fine-tuned models (see Appendix B in Yu et al. (2024)). We use the following Alpaca-based prompt (Taori et al., 2023) for training and evaluation:
@@ G.4 Matrix Factorization
Here, we provide the experimental details for Fig. 3. We compare Riemannian Gradient Descent (RGD) with the PoLAR parameterization (PO) to Gradient Descent (GD) with the Burer-Monteiro parameterization (BM). We set $\mathbf{A}\in\mathbb{R}^{m\times n}$ with $r_{A}=4$ evenly spaced singular values in the interval $1,\kappa]$ with $\kappa\in\{10,100\}$. For $\kappa=10$, we set $\eta=10^{-3}$ and $\gamma=1$ for PO + RGD and $\eta=10^{-3}$ for BM + GD. For $\kappa=100$, we use $\eta=10^{-4}$ for both methods. In the symmetric case, we use again $\eta=10^{-3}$ for $\kappa=10$ and $\eta=10^{-5}$ for $\kappa=100$.